OpenStudy (anonymous):

what is the simplified form of x-4/x^2-x-12 times x-4/x^2-8x+16

OpenStudy (anonymous):

\[\frac{x-4}{x^2-x-12}*\frac{x-4}{x^2-8x+16}= \frac{x-4}{(x-4)(x+3)}*\frac{x-4}{(x-4)^2}\]